### v1

# Charge radii measurements of exotic tin isotopes in the proximity of and

Abstract
We report nuclear charge radii for the isotopes 104-134Sn, measured using two different collinear laser spectroscopy techniques at ISOLDE-CERN. These measurements clarify the archlike trend in charge radii along the isotopic chain and reveal an odd-even staggering that is more pronounced near the =50 and =82 shell closures. The observed local trends are well described by both nuclear density functional theory and valence space in-medium similarity renormalization group calculations. Both theories predict appreciable contributions from beyond-mean-field correlations to the charge radii of the neutron-deficient tin isotopes. The models fall short, however, of reproducing the magnitude of the known transition probabilities, highlighting the remaining challenges in achieving a unified description of both ground-state properties and collective phenomena.
Introduction.– Understanding how nuclear shell structure and collective nuclear phenomena emerge from microscopic interactions between protons and neutrons is an essential step towards a unified description of the atomic nucleus. Collective nuclear properties are commonly enhanced in open-shell nuclei and significantly reduced around closed nuclear shells. With a closed proton shell at , and two predicted doubly magic isotopes, 100Sn and 132Sn, the tin isotopic chain provides a fertile testing ground for the development of theoretical ab-initio models [1, 2, 3, 4, 5, 6]. The properties of nuclei in the vicinity of these isotopes are expected to be accurately predicted by considering only a few valence nucleons. While the doubly magic nature of 132Sn has been well-established experimentally [7, 8, 9, 10, 11], significantly less data are available to understand the evolution of nuclear structure properties in the region from mid-shell down towards 100Sn. This is due to the considerable experimental challenges associated with producing and studying these neutron-deficient isotopes. Contradictory experimental evidence has been reported about the evolution of single-particle properties approaching 100Sn [12, 13, 14], and several questions remain open in connection with the robustness of the shell closures [1, 14, 15, 16, 17, 18, 19]. While the largest Gamow-Teller strength among all allowed -decays was found in 100Sn and supports a dominant single-particle behavior [1], the existence of low-lying energy levels and relatively large values, known down to 104Sn [15, 20, 21, 22], are indicative of a collective picture, in contrast with recent laser spectroscopy results on 49In [23] and 47Ag [16] isotopes.
In this work, we present nuclear charge radii measurements for the tin isotopes 104-134Sn, using two collinear laser spectroscopy (CLS) techniques [24]. The neutron-deficient isotopes, down to 104Sn, were investigated with the Collinear Resonance Ionization Spectroscopy (CRIS) setup [25], while the neutron-rich isotopes, up to 134Sn, were measured using the Collinear Laser Spectroscopy (COLLAPS) experiment [9]. Our measurement campaigns thus extend the available data [26, 27, 28] by four isotopes on the neutron-deficient side 104-107Sn, while completing the earlier reported radii in [9] with the odd-even isotopes including 133Sn. (We note that Fig. 7 of Ref. [29] contains a data point for the charge radius of 133Sn. This has been a tentative estimate, not based on actual measurement, and was unfortunately not removed from the final version of the paper.) This exceptionally long dataset, spanning thirty-one isotopes, serves to reveal intricate changes in the mean-square charge radii towards the neutron shell closures, which can be explored by nuclear density functional theory (DFT) [30, 31] and ab initio valence space in-medium similarity renormalization group (VS-IMSRG) calculations [32, 33].
Experiment.– Two CLS methods were employed to measure the charge radii of tin isotopes. Neutron-deficient isotopes, from 124Sn down to 104Sn, were measured at CRIS using two recently developed ionization schemes for probing the two atomic transitions (284 nm) and (281 nm) [34]. Neutron-rich isotopes, from 108Sn up to 134Sn were investigated at COLLAPS using two complementary transitions, (452.5 nm) and (286.3 nm). These transitions were chosen to maximize sensitivity to nuclear observables. Detailed information about the ion beam production and experimental technique used in the CRIS experiment, as well as the deduced isotope shifts and mean-square charge radii, can be found in the Supplemental Material. For the COLLAPS measurements, all details are described in [9, 35, 10]. The changes in the mean-square nuclear charge radii were extracted from the measured isotope shifts following the relation
where and are, respectively, the field-shift factor and the mass-shift factor of the transition , while and are the atomic masses of different isotopes.
| (fm2) | (fm) | ||
| 104* | 0 | (5) [41] | (5) |
| 105* | 5/2+ | (4) [38] | (4) |
| 106* | 0 | (3) [25] | (3) |
| 107* | 5/2+ | (3) [24] | (3) |
| 108 | 0 | (3) [16] | (2) |
| 109 | 5/2+ | (1) [16] | (2) |
| 110 | 0 | (3) [10] | (2) |
| 111 | 7/2+ | (1) [12] | (2) |
| 112 | 0 | (1) [6] | (1) |
| 113 | 1/2+ | (2) [7] | (1) |
| 114 | 0 | (1) [9] | (1) |
| 115 | 1/2+ | (1) [5] | (1) |
| 116 | 0 | (1) [7] | (1) |
| 117 | 1/2+ | (1) [4] | (1) |
| 118 | 0 | (1) [3] | (1) |
| 119 | 1/2+ | (1) [5] | (1) |
| 120 | 0 | (1) [2] | (1) |
| 121 | 3/2+ | (1) [5] | (1) |
| 122 | 0 | (1) [1] | (1) |
| 123 | 11/2- | (1) [3] | (1) |
| 124 | 0 | 0 | |
| 125 | 11/2- | (1) [4] | (1) |
| 126 | 0 | (1) [4] | (1) |
| 127 | 11/2- | (1) [8] | (1) |
| 128 | 0 | (1) [9] | (1) |
| 129 | 3/2+ | (2) [18] | (2) |
| 130 | 0 | (1) [15] | (2) |
| 131 | 3/2+ | (1) [25] | (3) |
| 132 | 0 | (1) [21] | (2) |
| 133* | 7/2- | (1) [12] | (2) |
| 134 | 0 | (2) [5] | (1) |
The atomic factors, and , were extracted from a standard King-plot analysis (as described in [9]) using muonic data as well as factors derived from elastic electron scattering [36]. Only the measured isotope shifts of the even stable isotopes were used in the King-plot analysis. Changes in charge radii extracted from the four transitions agree well with each other. The weighted averages were used to determine the nuclear charge radii of the tin isotopes based on the 124Sn reference value of fm of [36]. The results are presented in Table 1. The uncertainties in the extraction of charge radii are dominated by the atomic factor uncertainty.
Theory.– The measured radii are confronted with predictions utilizing both the VS-IMSRG and DFT methods. The VS-IMSRG results employ two interactions derived from chiral effective field theory, 1.8/2.0 (EM) [37, 38] and N2LOGO(394) [39]. The 1.8/2.0 (EM) interaction was fitted to data of few-nucleon systems and globally reproduces ground-state energies up to nuclei [40]. This interaction has been shown to provide a reasonably good description of local variations of nuclear charge radii in the medium-mass region [41], while underestimating their absolute values [42]. In contrast, the N2LOGO(394) interaction includes the -isobar degree of freedom and was fitted to few-body data, selected nuclei, and properties of infinite nuclear matter. This interaction improves the simultaneous reproduction of energies and radii [39]. Further details of the VS-IMSRG calculations are provided in the Supplemental Material.
The DFT-Fy(IVP) results were obtained with the HFB solver SkyAx for axially symmetric deformed configurations [43]. The computation of broken-pair states in odd- isotopes involves the breaking of time-reversal symmetry [44]. The applied Fy(IVP) functional [23] has been calibrated using the large dataset of ground-state properties of semi-magic nuclei from [45] augmented by the data on differential charge radii of the calcium [46], tin () and lead isotopic chains [23]. The calculated charge radii were obtained directly from the charge form factors containing relativistic corrections (including spin-orbit term) and contributions from nucleonic charge form factors [47]. The calibration of the functional produces statistical uncertainties from propagation of the uncertainties of the model parameters [45, 48]. Systematic uncertainties come from approximate angular momentum projection and from collective ground-state correlations stemming from low-lying quadrupole excitations [49]. The latter generally increase the radii and so introduce asymmetric uncertainties.
Discussion.– In Figure 1, we present a detailed comparison between our measured charge radii and predictions using both VS-IMSRG and DFT approaches using the experimentally determined spins. Figure 1(a) reveals that both interactions used in the VS-IMSRG calculations underestimate the absolute charge radii. This underestimation is primarily attributed to an insufficient description of saturation [50]. In the description of saturation in VS-IMSRG, three-nucleon (3N) forces play an important role. These cannot be fixed by information from NN scattering but largely cancel in the differential radii that can thus serve as a crucial test for VS-IMSRG calculations. The Fy(IVP) DFT model provides an improved description of the radii trend and the absolute radii compared to its former variant Fy(,HFB) employed in [9] for even-even tin isotopes, particularly in the neutron-rich region and across the kink at the shell-closure.
In Fig. 1(b) differential mean-square charge radii relative to 124Sn are shown. They show an archlike trend towards the known kink at the 132Sn doubly-magic shell closure. This general behavior superimposed on a linear increase is overall well described by both the DFT and VS-IMSRG calculations. The good agreement with a spherical Hartree-Fock (HF) calculation using the equal-filling approximation (labeled by “1.8/2.0 (EM) sph-HF” and discussed below) indicates that this is primarily driven by the bulk radial increase, whereas the correlations included in the VS-IMSRG calculations become important below . The collective ground-state correlations become sizable in the DFT calculations in the same region, and they move the result closer to experiment (as indicated with a light yellow band).
Finally, we study the odd-even staggering in Fig. 1(c). The addition of individual neutrons to the nucleus results in appreciable changes to its charge distribution due to pairing and deformation effects [41]. In our data, we observe an increased staggering as we approach the and shell closures. Notably, the characteristic inversion at , previously unobserved in tin, is now evident. The VS-IMSRG calculations for show improved agreement near these shell closures, although they tend to underestimate experiment in the mid-shell where a greater collectivity is expected. The Fy(IVP) calculations accurately predict the inversion across but systematically overestimate the staggering throughout the entire chain due to a slightly too large proton pairing predicted in mid-shell tin isotopes. It is to be noted that the estimated uncertainty in the radii below in Fy(IVP) is greater than the staggering magnitude; therefore, the results for these isotopes are omitted in Fig. 1(c).
To gauge the impact of cross-shell excitation [51] across , we extended our 1.8/2.0(EM) VS-IMSRG calculations to a valence space based on the 80Zr core. This model space did, however, not include the orbit, which is known to be crucial for capturing the nuclear structure effects for isotopes with [52]. Therefore, calculations with a 80Zr core were performed from up to only. The two choices of core should yield identical results if operators induced during the IMSRG flow were not truncated, so the discrepancy between them gives an indication of the VS-IMSRG truncation error. As shown in Fig. 1, the charge radii of the light tin isotopes predicted in VS-IMSRG show relatively low sensitivity to the assumed core. We therefore conclude the charge radii description can be captured by using a 100Sn core, with calculations performed in the range from to . (As we discuss below, this is not the case for observables).
We now focus on the even-even isotopes, which are simpler to interpret as the odd-neutron polarization is absent. The observed archlike trend in radii between shell closures is highlighted by removing the linear trend between the expected closed-shell isotopes 132Sn and 100Sn. The charge radius of 132Sn at the shell-closure is known. For 100Sn, the charge radius is estimated by fitting a parabolic function to the even-even radii down to 104Sn, following the same procedure described in Ref. [23]. In Fig. 2(a), the resulting archlike trend is shown after subtracting a bulk linear increase intersecting the charge radii at and .
This archlike behavior can be compared to the known data on the reduced transition probabilities in Fig. 2(b). These values provide insight into quadrupole collectivity along the isotopic chain. While the measured radii reveal a clearly reduced trend towards , the trend for the values is less evident due to the large systematic uncertainty of the experimental values for the neutron-deficient isotopes. The values computed with Fy(IVP) were deduced from collective correlations, using the same approach as the correlation correction to the charge radii whereby the BCS approximation was used for simplicity. It is to be noted that these values represent the sum over all transitions. Since the lowest transition usually accounts for 90% of the total strength, the Fy(IVP) values in Fig. 2(b) serve as upper estimates and should be reduced by 10%. The Fy(IVP) results generally yield an archlike trend with a pronounced maximum at mid-shell and an only moderate reduction toward neutron-deficient isotopes. However, the mid-shell peak in is exaggerated in a similar fashion as the collective corrections for the radii. This is due to an unusually small neutron pairing predicted in that region and thus calls for further model developments.
A similar archlike behavior of radii is obtained in the VS-IMSRG calculations, see Fig. 2(a). To explore the origin of this trend, spherical Hartree-Fock calculations were performed in the equal filling approximation [62], in which all -substates are filled equally, following the neutron orbitals in the order and . Hence, the observed trend is consistent with the charge radius following the neutron radius. The latter is mainly given by the sum of the single-particle mean-square radii weighted by their orbital occupation. The effect of configuration mixing is thus visualized by the comparison with the spherical HF and full VS-IMSRG results. The spherical HF results are not sensitive to collective effects, yet these calculations predict an overall archlike trend in radii. We conclude that – within the VS-IMSRG framework – the evolution of charge radii between the shell closures is dominated by the uniform isotropic growth of the mean field seen by the protons and is not due to many-body correlations. This result emphasizes that care should be taken when studying the correlations between quadrupole collectivity and nuclear charge radii [63, 64, 65, 66], as it is non-trivial to decouple the origin of different contributions to charge radii.
Finally, we note that the VS-IMSRG approach significantly underestimates the magnitude of the values, as seen in Fig. 2(b). As shown earlier [67, 68] VS-IMSRG calculations based on spherical reference states typically underestimate absolute values due to missing particle-hole excitations which are needed to capture the full collective strength. The extended valence space using the 80Zr core shows an improvement, with a large increase for the magnitude of the predicted values with respect to the results using a 100Sn core, but still significantly underestimates the experimental values. This further highlights the importance of cross-shell excitations for describing the measured values of the neutron-deficient tin isotopes, as previously suggested by large-scale shell model calculations [3]. Evidently, the many-body dynamics driving the behavior of the charge radii in Fig. 2(a) is distinct from that driving the quadrupole collectivity in Fig. 2(b).
Conclusions.– We interpreted the laser spectroscopy measurements of charge radii for 104-134Sn by DFT and VS-IMSRG approaches. Both are able to describe the general behavior of the radii. While the general features are better predicted near the shell closures, improved theory is needed to reproduce the observed odd-even staggering in the mid-shell, where the significant quadrupole collectivity is expected. Both DFT and VS-IMSRG calculations predict appreciable contributions from beyond-mean-field correlations to the charge radii of the neutron-deficient tin isotopes.
As previously noted in the references [69, 41, 70, 71], reproducing the absolute magnitude of the charge radii and values in VS-IMSRG remains a challenge. While an accurate reproduction of absolute charge radii in ab initio calculations is linked to developing improved nuclear forces with correct saturation, the underestimation of values is related to the lack of collective effects in the spherical VS-IMSRG employed in this work.
Within the VS-IMSRG framework, we have shown that the observed archlike trend of charge radii between shell closures is dominated by an expansion of the nucleus due to a gradual filling of neutron shells. This is in contrast with the conclusions obtained for values and other observables [23], which require the inclusion of higher-order many-body correlations and/or non-spherical reference states. Anticipated charge radii and measurements reaching closer to 100Sn will be of great interest to complete our understanding of this frontier region of the nuclear chart.
Acknowledgements.
We thank the ISOLDE technical group for their professional assistance. Useful comments from Ante Ravlić are gratefully acknowledged. This work was supported by ERC Consolidator Grant No. 648381 (FNPMLS); STFC grants ST/L005794/1, ST/L005786/1, (ST/L005670/1, ST/P004423/1, and Ernest Rutherford Grant No. ST/L002868/1; Projects GOA 15/010 and C14/22/104 from KU Leuven; the FWO-Vlaanderen project G080022N (Belgium); the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under grants DE-SC0021176 and DE-SC0013365, DE-SC0018083 and DE-SC0023175 (Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing); the Max-Planck Society; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 279384907 – SFB 1245; the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 101020842); the German Federal Ministry for Education and Research under Contract No. 05P15RDCIA and 05P24RD4; the Helmholtz International Center for FAIR (HIC for FAIR) within the LOEWE program by the State of Hesse; the European Union Grant Agreement 654002 and 262010 (ENSAR); National Key R&D Program of China (Contract No: 2018YFA0404403); the National Natural Science Foundation of China (No:11875073); NSERC under grants SAPIN-2018-00027, RGPAS-2018-522453, SAPIN-2024-0003, and the Arthur B. McDonald Canadian Astroparticle Physics Research Institute. TRIUMF receives funding via a contribution through the Digital Research Alliance of Canada. Computations were performed with an allocation of computing resources on Cedar at WestGrid and Compute Canada. JK acknowledges support from a Feodor-Lynen postdoctoral research fellowship funded by the Alexander-von-Humboldt Foundation. We acknowledge the computing resources provided by MIT and the Texas A&M high-performance computing cluster. TO, NS, and YT acknowledge the support by “Program for Promoting Researches on the Supercomputer Fugaku” (JPMXP1020200105, hp220174, hp210165). TM and NS acknowledge the support by JST ERATO Grant No. JPMJER2304, Japan. PGR thanks the regional computing center of the university Erlangen (RRZE) for support. This work was performed under the auspices of the U.S. Department of Energy, Office of Nuclear Physics (DOE NP) by Oak Ridge National Laboratory under Contract No. DE-AC05-00OR22725. DLB is supported in part through grants from the Romanian Ministry of Research, Innovation and Digitalization under contracts No. PN 23 21 01 06 and ELI-RO-RDI-2024-007 (ELITE). The data analysis used NumPy [72], SciPy [73], Pandas [74, 75], and iMinuit [76, 77]. The figures were produced using Matplotlib [78].Appendix A Supplemental Material
VS-IMSRG calculations
The VS-IMSRG calculations [32, 33] start from a Hamiltonian with two- and three-nucleon interactions represented in a space of 15 harmonic-oscillator shells with frequency MeV and the sufficiently large truncation for 3N matrix elements [79]. In the current study, we employed two chiral interactions, 1.8/2.0 (EM) [37, 38] and N2LOGO(394) [39]. Combined with the ensemble normal ordering method [80] to capture effects of 3N forces between valence nucleons, we decouple a valence space Hamiltonian using the Magnus expansion detailed in Ref. [81], where operators are truncated at the two-body level, the IMSRG(2) approximation. The radius operators are consistently transformed, and the charge radii are computed as in Ref. [82].
In the current study, three different valence spaces were adopted: The one-major shell space (above a 80Zr core), the neutron orbits above a 100Sn core, and the neutron above a 114Sn core. The VS-IMSRG calculations were performed with the imsrg++ code [83], and the effective Hamiltonians for the 100Sn core space were diagonalized with KSHELL code [84]. Since the exact diagonalizations of the valence-space Hamiltonians within the -shell space are not tractable, these calculations required the use of the recently developed quasi-particle vacua for nuclear shell-model calculations QVSM [85].
Experimental methodology
For the CRIS experiment, beams of stable and radioactive tin isotopes were produced from the bombardment of a heated LaCx target (1500-2000 ∘C) with a 1.4 GeV proton beam at the ISOLDE-CERN facility. The nuclear reaction recoils escape the target material through diffusion and effusion and enter a heated transfer line to an ion source. There, the tin atoms were selectively ionized with the resonant ionization laser ion source (RILIS) [86], and electrostatically accelerated to 40 keV. The ion-beam is steered through the high-resolution mass separator (HRS) magnets for isotope separation (). The separated beam is then sent into a cooler buncher linear Paul trap (ISCOOL) [87]. The ion bunches are released with a bunch length of 5 , and transported to the CRIS beamline, described in detail elsewhere [25]. Subsequently the ions were neutralized with sodium vapour at 230-250 ∘C within a charge-exchange-cell (CEC). The fraction of neutral atoms within the beam varied between 10–30 %. Any remaining ions were removed by electrostatic deflectors after the CEC. The atoms entered the interaction region held at a high vacuum of mbar to reduce the probability of collisional ionization of the neutral atom beam. Within the interaction region, each atom bunch was collinearly overlapped and synchronized in time with a series of laser pulses at 100 Hz repetition rate, which step-wise excited and non-resonantly ionized the atomic bunch. The two atomic transitions probed for studying the isotope shifts are (284 nm) and (281 nm). The two transitions were studied using the three- and two-step ionization schemes, as described in Ref. [34, 88]. These transitions were scanned with a speed below 5 MHz/s across the resonance peak using an injection-seeded single-mode titanium sapphire laser, operating at a repetition rate of 1 kHz, with a 20 MHz bandwidth in the fundamental output. The resonantly ionized ions are deflected onto an ETP DM291 MagneTOFTM detector with a timing resolution of 0.4 ns for event-per-event ion counting.
Example spectra of the Doppler corrected resonances observed when probing the transition for the even-even () isotopes are presented in Fig. 3, revealing the isotopic shift which contains the information of the changes in the nuclear charge radii .
Throughout the experiment, the ion beam acceleration voltage was measured, allowing accurate Doppler-correction of the spectra. The wavelength of the scanned step was regularly measured using a WSU-2 HighFinesse wavemeter, which simultaneously recorded a reference wavelength from a stabilized laser diode (Toptica DLC DL PRO 780) locked to the Doppler-free hyperfine transition of 87Rb (TEM CoSy) for correcting drifts of the wavemeter’s interferometer. Furthermore, data taking on the isotopes of interest was interleaved throughout the experiment campaign with measurements of selected reference isotope 112Sn to ensure any additional drifts were taken into account when determining the isotope shifts.
The transition in tin was predominantly used throughout the experiment, allowing spectroscopy of 21 tin isotopes ranging from 104Sn to 124Sn. Complementary measurements were performed using the transition down to 106Sn. During the spectroscopy of 114,116Sn using the transition, the laser power of the first step was increased by a factor five for enhancing ionization efficiency while increasing the linewidth from 230 MHz to 350 MHz total FWHM. Reference measurements of both 112Sn and 120Sn were performed under these identical conditions. Resonances were fitted using a Lorentz profile with free linewidth, since no Gaussian contribution was observed, with an average linewidth of 230 MHz for all isotopes except those acquired using a higher laser power. As an example, the fitted spectra extracted using the transition of the even-even isotopes are presented in Fig. 3.
| CRIS | COLLAPS | |||||
| (MHz) | (MHz) | (MHz) | (MHz) | |||
| 104 | 0 | |||||
| 105 | 5/2+ | |||||
| 106 | 0 | |||||
| 107 | 5/2+ | |||||
| 108 | 0 | |||||
| 109 | 5/2+ | |||||
| 110 | 0 | |||||
| 111 | 7/2+ | |||||
| 112 | 0 | 0 | 0 | |||
| 113 | 1/2+ | |||||
| 114 | 0 | |||||
| 115 | 1/2+ | |||||
| 116 | 0 | |||||
| 117 | 1/2+ | |||||
| 118 | 0 | |||||
| 119 | 1/2+ | |||||
| 120 | 0 | |||||
| 121 | 3/2+ | |||||
| 122 | 0 | |||||
| 123 | 11/2- | |||||
| 124 | ] | |||||
| 125 | 11/2- | |||||
| 126 | 0 | ] | ||||
| 127 | 11/2- | |||||
| 128 | 0 | |||||
| 129 | 3/2+ | |||||
| 130 | ||||||
| 131 | 3/2+ | |||||
| 132 | ||||||
| 133 | 7/2- | |||||
| 134 | 0 | |||||
The determined isotope shifts relative to 124Sn are provided in Tab. 2 for both transitions. In order to obtain from the isotope shifts measured at CRIS relative to 112Sn, the following procedure was used. First, the isotope shifts relative to 124Sn were calculated and the statistical uncertainties of the isotope of interest and that of the new reference 124Sn were added in quadrature. To determine the total uncertainty, we combined the statistical uncertainty with the systematic uncertainty specific to each isotope. In this particular case, the systematic uncertainty was not derived from a potential voltage offset. Instead, it stemmed from laser drifts that could occur during the measurement of different isotopes, whether in positive or negative directions. For the isotope 112Sn, we used the systematic uncertainty associated with 124Sn. This total uncertainty was then used to perform the King plot for the individual transitions. For the COLLAPS data, only the statistical uncertainty was taken into account, since the systematic uncertainty is dominated by the scaling uncertainty of the voltage divider that transfers into a mass-shift like behavior in the isotopic chain and is thus captured by the King-plot procedure. The final value for as listed in Table 1 is then obtained as the weighted mean of the measurements of the four transitions. The final uncertainty has been chosen as the smallest uncertainty that was obtained from one specific transition, but no further reduction was considered due to the strong correlations via the respective field shift uncertainty. It was found that the size of this uncertainty covers approximately the observed range of variation of the results in the four transitions.
| Transition | 284 nm | 281 nm | 286 nm | 453 nm |
| (GHz/fm2) | 2.51 (53) | 2.68 (72) | 2.42 (47) | 2.74 (57) |
| (GHz u) | -228 (447) | -838 (610) | -492 (397) | -456 (480) |
| (GHz u) | 1888 (17) | 1428 (21) | 1543 (15) | 1850 (17) |
| ( fm2 u) | 843 | 846 | 841 | 842 |
The mass shift and field shift factors obtained in the King plots, which are required to obtain the values in the respective transitions, are summarized in Table LABEL:table:atomic_factors. The differential mean-square charge radii and their uncertainties for each transition individually can be obtained using the values in Tab. 2 and Tab. LABEL:table:atomic_factors, according to with the mass factor . In the evaluation of the uncertainty on , the optimized ensures that no correlations between and exist in the error propagation.